Ralph spent one quarter of his allowance at the school tuckshop purchasing lunch Monday and two thirds of the remainder at the bookstore on Wednesday. If he had $4.50 left in his wallet, how much is Ralph's weekly allowance?
Answer by Theo(13342)

x = allowance.
cost for lunch = x/4
remainder = 3x/4
cost for books = 2/3 * (3x/4) = 2x/4
remainder = 3x/4 - 2x/4 = x/4
remainder = 4.50 which leads to:
4.50 = x/4
multiply both sides by 4 to get:
x = 18.00
to confirm, go back to the original problem and work it through with x = 18.00
allowance = 18.00
cost for lunch = 18/4 = 4.5
remainder = 18.00 - 4.5 = 13.5
cost for books = 2/3 * remainder
cost for books = 2/3 * 13.5 = 27/3 = 9
remainder = 13.5 - 9 = 4.50
everything checks out.
allowance is $18.00